OpenStudy (yanasidlinskiy):

Multiply or divide. Solve the equation. 5x + 10/x+2 multiplied by 2x / 4x - 10 Solve the equation.

OpenStudy (anonymous):

Just multiply the across the numerators and multiply across the denominators. You will distribute the 2x across (5x+10) in your answer's numerator, and multiply the binomials (4x-10)(x+2) in you answer's denominator.

OpenStudy (anonymous):

\[\frac{a}{b}\times\frac{c}{d} = \frac{ac}{bd}\]

OpenStudy (yanasidlinskiy):

So would the answer be 7x+10/ 5x-8??